The New First Two Year Curriculum A Progress Report

Margaret M. Reek

Department of Computer Science, RIT

Abstract

The first year of the new lower division curriculum is almost completed, and plans are under way to implement the second year courses. The new curriculum represents a radical departure from the traditional ways of teaching computer science, both in terms of content, order of presentation of topics, and in the way the course materials are developed. This colloquium will update faculty on the philosophy, process and status of this development effort.

The overall curriculum philosophy of taking a top-down approach to teaching software concepts, introducing software engineering principles early and reinforcing them throughout the program will be discussed. The brief overview of the structure of and major topics covered in each class will be presented to show how the philosophy is translated into a course structure. The reactions of the faculty and students to the courses and philosophy will be examined.

The manner in which the course materials have been developed represent a major paradigm shift. The team based approach to curriculum and materials development has been central to our successes to date. The philosophy will be discussed, along with a brief overview of some of the tools used to facilitate these efforts.

Coffee and cookies will be served